A university law school accepts 3 out of every 8 applicants. If the school accepted 240 students find how many applications they received
Answer by stanbon(75887)

A university law school accepts 3 out of every 8 applicants. If the school accepted 240 students find how many applications they received
--------------
Ans: # = (3/8)240 = 3*30 = 90
--------------
Cheers,
Stan H.
-------------